Two integers have the property that the larger one is 6 less than 9 times the smaller one. Their sum is 64. What are the two integers?
Answer by emargo19(101)

Let the 1st number be x
Let the 2nd number be 9x-6
The sum= x+9x-6
x+9x-6=64
10x=70
x=7
Ist integer=7
2nd integer=9x-6
=9(7)-6
=63-6
=57
Ans= 7, 57